# Moments of inertia of rare-earth nuclei and the nuclear time-odd mean fields within exact solutions of the adiabatic theory

### v1

@@ Abstract
We systematically analyse the nuclear moments of inertia determined within the Skyrme and Gogny density-functional theories. The time-odd mean fields generated by collective rotation are self-consistently determined by an exact iterative solution of the adiabatic time-dependent Hartree-Fock-Bogoliubov (ATDHFB) equations. Although details of the results depend on the functional used, the calculated moments of inertia are in good overall agreement with the experimental data, with no adjustable parameters. To show the essential importance of the time-odd mean fields, the ATDHFB moments of inertia are compared with those obtained from the Inglis-Belyaev formula. For Skyrme density functionals, we find strong correlations between the effective mass and the impact of the time-odd mean fields on inertia’s rotational and vibrational moments.
@@
journal: Physics Letter B
Collective motion provides a deep insight into the nuclear reaction mechanism and the quantum dynamics of many-nucleon systems. Of particular interest is the low-energy collective motion, usually accompanied by large electromagnetic moments and transition rates. To fully understand them, we have to scrutinize the collective inertia, which measures the quantum system’s resistance to collective motion.
@@
Despite the inconsistency caused by neglecting the time-odd mean fields, the cranking approximation was used in the theoretical studies of the collective inertia, from the calculation of the low-energy spectra [8] to the evaluation of nuclear fission half-life [9]. Such an inconsistency led to the enhancement factor adopted to correct the moment of inertia obtained from the cranking approximation [10]. To alleviate the inconsistency, a method based on the expansion of the inertia matrix was developed to solve the ATDHFB equation in the absence of pairing, which is able to evaluate the rotational moment of inertia with a high accuracy of the order of 1% [11]. Recently, the similarity of ATDHFB method and the finite-amplitude method at zero frequency was realized, and a quasiparticle finite-amplitude method on top of the Skyrme energy density functional was developed to evaluate the Thouless-Valatin rotational moment of inertia [12]. With the pairing strength adjusted to the experimental pairing gaps, the rotational moment of inertia in several axially deformed nuclei from the rare-earth and heavy-actinide regions was well reproduced, and the large and mass-dependent deviations from the values evaluated by the cranking approximation was confirmed. For the full five-dimensional quadrupole adiabatic inertia, the finite-amplitude method was also used in Ref. [13].
@@
The one-body quasiparticle Hamiltonian $\mathcal{H}$ is the derivative of the energy of the nuclear system over the density. In the non-relativistic DFT it is usually determined using either the zero-range Skyrme  [16] or finite-range Gogny [17] functional.
@@
where $\mathcal{R}_{0}(q)$ and $\mathcal{H}_{0}(q)$ are, respectively, the static (time-even) density and Hamiltonian along the collective path parameterised by the collective variable $q$. The corresponding time-odd density $\mathcal{R}_{1}$ and Hamiltonian $\mathcal{H}_{1}$ are small corrections generated by the collective motion and $\dot{q}$ denotes the collective velocity. For the collective rotation, the velocity is proportional to the angular momentum operator [18]; for the collective vibration, it can be evaluated from numerical differentiation [14]. Since up to the first order $\mathcal{H}_{1}$ depends linearly on $\mathcal{R}_{1}$, the ATDHFB equation (2) is a linear equation for $\mathcal{R}_{1}$ of many dimensions, which requires an unmanageable inversion of the huge stability matrix [2].
@@
In summary, we proposed a novel iterative method to solve the ATDHFB equations exactly and we applied the method to investigate the moment of inertia in deformed nuclei. For two Skyrme functionals SkM* and UNEDF1 and one Gogny functional D1S, we determined the rotational moment of inertia of the even-even nuclei between gadolinium and osmium. We obtained a good overall agreement with experimental data, although improvements in functional parameterizations, especially in the pairing channel, are required. The efficiency of our method bids well for the possibility of including the rotational moments of inertia in future adjustments of nuclear functionals.

### v2

@@ Abstract
We systematically analyse the nuclear moments of inertia determined within the Skyrme and Gogny density functional theories. The time-odd mean fields generated by collective rotation are self-consistently determined by a novel exact iterative solution of the adiabatic time-dependent Hartree-Fock-Bogoliubov (ATDHFB) equations. Although details of the results depend on the functional used, the calculated moments of inertia are in good overall agreement with the experimental data, with no adjustable parameters. To show the essential importance of the time-odd mean fields, we compared the ATDHFB moments of inertia with those obtained from the Inglis-Belyaev formula. For Skyrme density functionals, we find strong correlations between the effective mass and the impact of the time-odd mean fields on the inertia’s rotational and vibrational moments.
@@
journal: Physics Letters B
Collective motion provides a deep insight into the nuclear reaction mechanism and the quantum dynamics of many-nucleon systems. The low-energy collective motion is particularly interesting, usually accompanied by large electromagnetic moments and transition rates. To fully understand them, we must scrutinise the collective inertia, which measures the quantum system’s resistance to collective motion.
@@
Despite the inconsistency caused by neglecting the time-odd mean fields, the cranking approximation was used in the theoretical studies of the collective inertia, from the calculation of the low-energy spectra [8] to the evaluation of nuclear fission half-life [9]. Such an inconsistency led to the enhancement factor adopted to correct the moment of inertia obtained from the cranking approximation [10]. To alleviate the inconsistency, a method based on the expansion of the inertia matrix was developed to solve the ATDHFB equation in the absence of pairing, which can evaluate the rotational moment of inertia with high accuracy of the order of 1% [11]. Recently, the similarity of the ATDHFB method and the finite-amplitude method at zero frequency was realised, and a quasiparticle finite-amplitude method on top of the Skyrme energy density functional was developed to evaluate the Thouless-Valatin rotational moment of inertia [12]. With the pairing strength adjusted to the experimental pairing gaps, the rotational moment of inertia in several axially deformed nuclei from the rare-earth and heavy-actinide regions was well reproduced, and the large and mass-dependent deviations from the values evaluated by the cranking approximation was confirmed. The finite-amplitude method was also used in Ref. [13] for the full five-dimensional quadrupole adiabatic inertia.
@@
The one-body quasiparticle Hamiltonian $\mathcal{H}$ is the derivative of the energy of the nuclear system over the density. In the non-relativistic DFT, it is usually determined using either the zero-range Skyrme  [16] or finite-range Gogny [17] functional.
@@
where $\mathcal{R}_{0}(q)$ and $\mathcal{H}_{0}(q)$ are, respectively, the static (time-even) density and Hamiltonian along the collective path parameterised by the collective variable $q$. The corresponding time-odd density $\mathcal{R}_{1}$ and Hamiltonian $\mathcal{H}_{1}$ are small corrections generated by the collective motion, and $\dot{q}$ denotes the collective velocity. For the collective rotation, the velocity is proportional to the angular momentum operator [18]; for the collective vibration, it can be evaluated from numerical differentiation [14]. Since up to the first order $\mathcal{H}_{1}$ depends linearly on $\mathcal{R}_{1}$, the ATDHFB equation (2) is a linear equation for $\mathcal{R}_{1}$ of many dimensions, which requires an unmanageable inversion of the huge stability matrix [2].
@@
In summary, we proposed a novel iterative method to solve the ATDHFB equations exactly and we applied it to investigate the moment of inertia in deformed nuclei. For two Skyrme functionals SkM* and UNEDF1 and one Gogny functional D1S, we determined the rotational moment of inertia of the even-even nuclei between gadolinium and osmium. We obtained a good overall agreement with experimental data, although improvements in functional parameterizations, especially in the pairing channel, are required. The efficiency of our method bids well for the possibility of including the rotational moments of inertia in future adjustments of nuclear functionals.